OpenStudy (anonymous):

Simple question - Need to know if I put this in the correct FORM READ BELOW FOR EQUATION

OpenStudy (anonymous):

\[x ^{2}+4y ^{2}+2y=6x+3y ^{2}-1\] IS A CIRCLE Vertex (3,-1) Radius (9) Correct form: \[(x-3)^{2}-\frac{ y ^{2} }{3}-\frac{ y }{ 6 }+\frac{ y ^{2} }{ 4 }=\frac{ -2 }{ 3 } \] Is this right?

OpenStudy (anonymous):

@satellite73 please take a look real quick if you can

geerky42 (geerky42):

Not really correct form. you need to get y to look like what x look like, for example: \[(x-h)^2 + (y-k)^2\]

OpenStudy (anonymous):

How do I do that? I've never done a problem like this before Can I combine the y^2 together?

OpenStudy (anonymous):

@Luigi0210

geerky42 (geerky42):

yeah you can.

geerky42 (geerky42):

\(-\dfrac{y^2}{3} + \dfrac{y^2}{4}\)

OpenStudy (anonymous):

Doesn't the denominator have to be the same? 12?

OpenStudy (anonymous):

@mathslover @phi @texaschic101 I need help on this please!!I have class in 40 min!!

mathslover (mathslover):

Yes, the denominator has to be same in order to add : LCM of 3 and 4 is 12 so, \(\dfrac{-4y^2 + 3y^2}{12}\) = \(\dfrac{-y^2}{12}\)

OpenStudy (anonymous):

so then, it would be \[(x-3)^{?}-\frac{ y ^{?} }{ 12 }+\frac{ y }{ 6 }=\frac{ -2 }{ }\]

OpenStudy (anonymous):

* -2/3

OpenStudy (phi):

based on the info Vertex (3,-1) Radius (9) the equation should transform to \[ (x-3)^2 + (y+1)^2 = 81 \]

OpenStudy (phi):

though I assume you mean center (not vertex)

OpenStudy (anonymous):

Yes Center is (3,1)

OpenStudy (anonymous):

(3,-1)****** So (x-3)^2+(y+1)^2=81 still is correcr?

OpenStudy (phi):

I get a radius of 3 (not 9)

mathslover (mathslover):

Well the given equation can be written as : \(x^2 + y^2 + 2y -6x + 1 = 0 \) \((x^2 - 6x) + (y^2 + 2y + 1) =0 \) \((x^2 -6x + 9 -9) + (y+1)^2 = 0\) \((x-3)^2 + (y+1)^2 = 9\)

OpenStudy (phi):

\[ x ^{2}+4y ^{2}+2y=6x+3y ^{2}-1 \\ x^2 -6x + y^2 +2y = -1 \\ \] complete the square for x by add (-6/2)^2 to both sides complete the square for y by adding (2/2)^2 to both sides

mathslover (mathslover):

But, I am not getting 81 in RHS... @phi - I think, the question is a little bit wrong, the radius should be 3..

OpenStudy (phi):

yes, I think mighty figured out the wrong radius.

mathslover (mathslover):

Yeah...

OpenStudy (anonymous):

Yes, my fault. It's the SQUARE ROOT of 9, not 9. So (x-3)^2+(y+1)^2=9 is correct

mathslover (mathslover):

yeah

OpenStudy (anonymous):

Man, thank you guys so much

mathslover (mathslover):

Thanks to phi :)

OpenStudy (phi):

yes. If you complete the square (see above for the details posted by math) you get that equation

OpenStudy (anonymous):

mathslover, if you're still there I have a quick Q for ya!

mathslover (mathslover):

Yes, sure! I am always here.

OpenStudy (anonymous):

in the equation (x-2)^2 OVER 9 + (y-1)^2 OVER 4 =1 (an Ellipse) is the center (2,1)?

OpenStudy (anonymous):

@mathslover

mathslover (mathslover):

Not studied Ellipse yet :( Very sorry.. @amistre64 may help you!

OpenStudy (amistre64):

ellipses are just flattened circles

OpenStudy (anonymous):

Nevermind, sorry I got it! It's (2,-1). I put (y-1) instead of (Y+1)

OpenStudy (amistre64):

\[\frac{x^2}{a^2}+\frac{y^2}{b^2}=1\] if a=b, its a circle, otherwise its an ellipse if one oof the terms iis negatiivee iits a hyperbola, if only one term is squared, it a parrabola

OpenStudy (amistre64):

im thinking my keypad is getting old ...

OpenStudy (anonymous):

Haha I know the feeling. OpenStudy is glitching up for me today. Thanks for the help, I understand it now!

OpenStudy (amistre64):

good luck :)